# Fast Benchmarking of Asynchronous Multi-Fidelity Optimization on Zero-Cost Benchmarks

### v1

@@ 1  Introduction
Figure 1: The simplest codeblock example of how our wrapper works. Left: a codeblock example without our wrapper (naïve simulation). We let each worker call sleep for the time specified by the queried result. This implementation is commonly used to guarantee correctness, as research often requires running optimizers from other researchers. Right: a codeblock example with our wrapper (multi-core simulation). Users only need to wrap the objective function with our module and remove the line for sleeping. In the end, both codeblocks yield identical results.
@@ 2  Background
In this section, we define our problem setup. Throughout the paper, we assume minimization problems of an objective function  [^3] As mentioned in Appendix B, we can also simulate with multi-objective optimization and constrained optimization. $f(\bm{x}):\mathcal{X}\rightarrow\mathbb{R}$ defined on the search space $\mathcal{X}\coloneqq\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{D}$ where $\mathcal{X}_{d}\subseteq\mathbb{R}$ is the domain of the $d$-th HP. Furthermore, we define the (predictive) actual runtime function $\tau(\bm{x}):\mathcal{X}\rightarrow\mathbb{R}_{+}$ of the objective function given an HP configuration $\bm{x}$. Although $f(\bm{x})$ and $\tau(\bm{x})$ could involve randomness, we only describe the deterministic version for the notational simplicity. In this paper, we use $\bm{x}^{(n)}$ for the $n$-th sample and $\bm{x}_{n}$ for the $n$-th observation and we would like to note that they are different notations. In asynchronous optimization, the sampling order is not necessarily the observation order, as certain evaluations can take longer. For example, if we have two workers and the runtime for the first two samples are $\tau(\bm{x}^{(1)})=200$ and $\tau(\bm{x}^{(2)})=100$, $\bm{x}^{(2)}$ will be observed first, yielding $\bm{x}_{1}=\bm{x}^{(2)}$ and $\bm{x}_{2}=\bm{x}^{(1)}$.
@@ 2.1  Asynchronous Optimization on Zero-Cost Benchmarks
The problems of this setting are that (1) the policy is conditioned on $\mathcal{D}_{N}$ and the order of the observations must be preserved and (2) each worker must wait for the other workers to match the order to be realistic. While an obvious approach is to let each worker wait for the queried runtime $\tau^{(n)}$ as in Figure 1 (Left), it is a waste of energy and time. To address this problem, we need a wrapper as in Figure 1 (Right).
@@ 2.2  Related Work
Although there have been many HPO benchmarks invented for MFO such as HPOBench (Eggensperger et al.,, 2021), NASLib (Mehta et al.,, 2022), and JAHS-Bench-201 (Bansal et al.,, 2022), none of them provides a module to allow researchers to simulate runtime internally. Other than HPO benchmarks, many HPO frameworks handling MFO have also been developed so far such as Optuna (Akiba et al.,, 2019)), SMAC3 (Lindauer et al.,, 2022), Dragonfly (Kandasamy et al.,, 2020), and RayTune (Liaw et al.,, 2018). However, no framework above considers the simulation of runtime. Although HyperTune (Li et al.,, 2022) and SyneTune (Salinas et al.,, 2022) are internally simulating the runtime, we cannot simulate optimizers of interest if the optimizers are not introduced in the packages. This restricts researchers in simulating new methods, hindering experimentation and fair comparison. Furthermore, their simulation backend assumes that optimizers take the ask-and-tell interface and it requires the reimplementation of optimizers of interest in their codebase. Since reimplementation is time-consuming and does not guarantee its correctness without tests, it is helpful to have an easy-to-use Python wrapper around existing codes. Note that this work extends previous work (Watanabe, 2023a, ), by adding the handling of optimizers with non-negligible overhead and the empirical verification of the simulation algorithm.
@@ 3  Automatic Waiting Time Scheduling Wrapper
As an objective function may take a random seed and fidelity parameters in practice, we denote a set of the arguments for the $n$-th query as $\bm{a}^{(n)}$. In this section, a job means to allocate the $n$-th queried HP configuration $\bm{x}^{(n)}$ to a free worker and obtain its result $r^{(n)}\coloneqq(f^{(n)},\tau^{(n)})=(f(\bm{x}^{(n)}|\bm{a}^{(n)}),\tau(\bm{x}^{(n)}|\bm{a}^{(n)}))$. Besides that, we denote the $n$-th chronologically ordered result as $r_{n}$. Our wrapper is required to satisfy the following features:
@@
Note that an example of the restart of evaluation could be when we evaluate DL model instantiated with HP $\bm{x}$ for $20$ epochs and if we want to then evaluate the same HP configuration $\bm{x}$ for $100$ epochs, we start the training of this model from the $21$st epoch instead of from scratch using the intermediate state. To achieve these features, we chose to share the required information via the file system and create the following JSON files that map:
@@
Algorithm 1 outlines the pseudocode.
We additionally provide an approach that also extends to the ask-and-tell interface by providing a Single-Core Simulator (SCS) for single-core scenarios (details omitted for brevity). While the Multi-Core Simulator (MCS) wraps optimizers running $P$ cores or $P$ workers, SCS runs only on a single core and simulates a $P$-worker run. Unlike previous work (Watanabe, 2023a, ), Algorithm 1 handles expensive optimizers by checking individual worker wait times. However, this check complicates race conditions, making it hard to guarantee the correctness of implementation. For this reason, empirical verification through edge cases is provided in the next section.
@@
4:         $\triangleright$ Condition 1: The new fidelity input must have higher fidelity
5:         $\triangleright$ Condition 2: The registration of $s^{(N+1)}$ must have happened before the sample
@@
7:     Query the result $(f^{(N+1)},\tau^{(N+1)})$
8:     Calibrate runtime $\tau^{(N+1)}\leftarrow\tau^{(N+1)}-\tau$
@@
11:     $\triangleright$ $k$ is the number of results from the other workers appended during the wait
@@
16:$\pi$ (an optimizer policy), get_n_results (a function that returns the number of recognized results by our wrapper. The results include the ones that have not been recognized by the optimizer yet.).
@@ 4  Empirical Algorithm Verification on Test Cases
In this section, we verify our algorithm using some edge cases. Throughout this section, we use the number of workers $P=4$. We also note that our wrapper behavior depends only on returned runtime at each iteration in a non-continual setup and it is sufficient to consider only runtime $\tau^{(n)}$ and sampling time $t^{(n)}$ at each iteration. Therefore, we use a so-called fixed-configuration sampler, which defines a sequence of HP configurations and their corresponding runtimes at the beginning and samples from the fixed sequence iteratively. More formally, assume we would like to evaluate $N$ HP configurations, then the sampler first generates $\{\tau^{(n)}\}_{n=1}^{N}$ and one of the free workers receives an HP configuration at the $n$-th sampling that leads to the runtime of $\tau^{(n)}$. Furthermore, we use two different optimizers to simulate the sampling cost:
@@
    Expensive Optimizer: that sleeps for $c(|\mathcal{D}|+1)$ seconds before giving $\tau^{(n)}$ to a worker where $|\mathcal{D}|$ is the size of a set of observations and $c\in\mathbb{R}_{+}$ is a proportionality constant, and
@@
    Cheap Optimizer: that gives $\tau^{(n)}$ to a worker immediately without sleeping.
@@ 4.1  Quantitative Verification on Random Test Cases
(a) Cheap optimizer
(a) Cheap optimizer
(b) Expensive optimizer with $c=5\times 10^{-2}$
Figure 3: The return order verification results. When we use our wrapper, the red dots are obtained. If all the dots are aligned on $y=x$, it implies that the return order in a simulation with our wrapper and that in its naïve simulation perfectly match. As expected, the red dots completely overlap with $y=x$. See the text for the plot details.
We test our algorithm quantitatively using some test cases. The test cases $\{\tau^{(n)}\}_{n=1}^{N}$ where $N=100$ for this verification were generated from the following distributions: 1. Uniform $\frac{T}{c}\sim\mathcal{U}(0,2)$, 2. Exponential $\frac{T}{c}\sim\mathrm{Exp}(1)$, 3. Pareto $\frac{T+1}{c}\sim\mathcal{P}(\alpha=1)$, and 4. LogNormal $\ln\frac{\sqrt{e}T}{c}\sim\mathcal{N}(0,1)$, where $T$ is the probability variable of the runtime $\tau$ and we used $c=5$. Each distribution uses the default setups of numpy.random and we calibrated each distribution except for the Pareto distribution so that the expectation becomes $5$. Furthermore, we used the cheap optimizer and the expensive optimizer with $c=5\times 10^{-2}$. As $N=100$, the worst sampling duration for an expensive optimizer will be $5$ seconds. As we can expect longer waiting times for the expensive optimizer, it is more challenging to yield the precise return order and the precise simulated runtime. Hence, these test cases empirically verify our implementation if our wrapper passes every test case.
@@
We performed the following procedures to check whether the obtained return orders are correct: (1) run optimizations with the naïve simulation (NS), i.e. Figure 1 (Left) and without our wrapper, i.e. Figure 1 (Right), (2) define the trajectories for each optimization $\{\tau^{\mathrm{NS}}_{n}\}_{n=1}^{N}$ and $\{\tau_{n}\}_{n=1}^{N}$, (3) sort $\{\tau_{n}\}_{n=1}^{N}$ so that $\{\tau_{i_{n}}\}_{n=1}^{N}=\{\tau^{\mathrm{NS}}_{n}\}_{n=1}^{N}$ holds, and (4) plot $\{(n,i_{n})\}_{n=1}^{N}$ (see Figure 3). If the simulated return order is correct, the plot $\{(n,i_{n})\}_{n=1}^{N}$ will look like $y=x$, i.e. $(n,n)$ for all $n\in[N]$, and we expect to have such plots for all the experiments. For comparison, we also collect $\{\tau_{n}\}_{n=1}^{N}$ without our wrapper, i.e. Figure 1 (Left) without time.sleep in Line 4. As seen in Figure 3, our wrapper successfully replicates the results obtained by the naïve simulation. The test cases by the Pareto distribution are edge cases because it has a heavy tail and it sometimes generates configurations with very long runtime, leading to blue dots located slightly above the red dots. Although this confuses the implementation without our wrapper completely, our wrapper appropriately handles the edge cases.
(a) Cheap optimizer
@@
We check whether the simulated runtimes at each iteration were correctly calculated using the same setups. Figure 4 presents the simulated runtimes for each setup. As can be seen in the figures, our wrapper got a relative error of $1.0\times 10^{-5}\sim 1.0\times 10^{-3}$. Since the expectation of runtime is $5$ seconds except for the Pareto distribution, the error was approximately $0.05\sim 5$ milliseconds and this value comes from the communication overhead in our wrapper before each sampling. Although the error is sufficiently small, the relative error becomes much smaller when we use more expensive benchmarks that will give a large runtime $\tau^{(n)}$.
@@ 4.2  Performance Verification on Actual Runtime Reduction
In the previous sections, we verified the correctness of our algorithms and empirically validated our algorithms. In this section, we demonstrate the runtime reduction effect achieved by our wrapper. To test the runtime reduction, we optimized the multi-fidelity 6D Hartmann function 444 We set the runtime function so that the maximum runtime for one evaluation becomes 1 hour. More precisely, we used $10\times r(\bm{z})$ instead of $r(\bm{z})$ in Appendix A.2.  (Kandasamy et al.,, 2017) using random search with $P=4$ workers over 10 different random seeds. In the noisy case, we added a random noise to the objective function. We used both MCS and SCS in this experiment and the naïve simulation. Figure 5 (Left) shows that both MCS and SCS perfectly reproduced the results by the naïve simulation while they finished the experiments $6.7\times 10^{3}$ times and $1.3\times 10^{5}$ times faster, respectively. Note that it is hard to see, but the rightmost curve of Figure 5 (Left) has the three lines: (1) Simulated Runtime (MCS), (2) Simulated Runtime (SCS), and (3) Actual Runtime (Naïve), and they completely overlap with each other. SCS is much quicker than MCS because it does not require communication between each worker via the file system. Although MCS could reproduce the results by the naïve simulation even for the noisy case, SCS failed to reproduce the results because the naïve simulation relies on multi-core optimization, while SCS does not use multi-core optimization. This difference affects the random seed effect on the optimizations. However, since SCS still reproduces the results for the deterministic case, it verifies our implementation of SCS. From the results, we can conclude that while SCS is generally quicker because it does not require communication via the file system, it may fail to reproduce the random seed effect since SCS wraps an optimizer by relying on the ask-and-tell interface instead of using the multi-core implementation provided by the optimizer.

### v3

@@ 1  Introduction
Figure 1: The simplest codeblock example of how our wrapper works. Left: a codeblock example without our wrapper (naïve simulation). We let each worker call sleep for the time specified by the queried result. This implementation is commonly used to guarantee correctness, as research often requires us to run optimizers from other researchers. Right: a codeblock example with our wrapper (multi-core simulation). Users only need to wrap the objective function with our module and remove the line for sleeping. In the end, both codeblocks yield identical results.
@@ 2  Background
In this section, we define our problem setup. Throughout the paper, we assume minimization problems of an objective function  [^3] As mentioned in Appendix B, we can also simulate with multi-objective optimization and constrained optimization. $f(\bm{x}):\mathcal{X}\rightarrow\mathbb{R}$ defined on the search space $\mathcal{X}\coloneqq\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{D}$ where $\mathcal{X}_{d}\subseteq\mathbb{R}$ is the domain of the $d$-th HP. Furthermore, we define the (predictive) actual runtime function $\tau(\bm{x}):\mathcal{X}\rightarrow\mathbb{R}_{+}$ of the objective function given an HP configuration $\bm{x}$. Although $f(\bm{x})$ and $\tau(\bm{x})$ could involve randomness, we only describe the deterministic version for the notational simplicity. In this paper, we use $\bm{x}^{(n)}$ for the $n$-th sample and $\bm{x}_{n}$ for the $n$-th observation and we would like to note that they are different notations. In asynchronous optimization, the sampling order is not necessarily the observation order, as certain evaluations can take longer. For example, if we have two workers and the runtime for the first two samples are $\tau(\bm{x}^{(1)})=200$ and $\tau(\bm{x}^{(2)})=100$, $f(\bm{x}^{(2)})$ will be observed first, yielding $\bm{x}_{1}=\bm{x}^{(2)}$ and $\bm{x}_{2}=\bm{x}^{(1)}$.
@@ 2.1  Asynchronous Optimization on Zero-Cost Benchmarks
The problems of this setting are that (1) the policy $\pi$ is conditioned on $\mathcal{D}_{N}$, which is why the order of the observations must be preserved, and (2) each worker must wait for the other workers to match the order to be realistic. While an obvious approach is to let each worker wait for the queried runtime $\tau^{(n)}$ as in Figure 1 (Left), it is a waste of energy and time. To address this problem, we need a wrapper as in Figure 1 (Right).
@@ 2.2  Related Work
Although there have been many HPO benchmarks invented for MFO such as HPOBench (Eggensperger et al.,, 2021), NASLib (Mehta et al.,, 2022), and JAHS-Bench-201 (Bansal et al.,, 2022), none of them provides a module to allow researchers to simulate runtime internally. We defer the survey by Li and Li, (2024) for the details of MFO. Other than HPO benchmarks, many HPO frameworks handling MFO have also been developed so far such as Optuna (Akiba et al.,, 2019)), SMAC3 (Lindauer et al.,, 2022), Dragonfly (Kandasamy et al.,, 2020), and RayTune (Liaw et al.,, 2018). However, no framework above considers the simulation of runtime. Although HyperTune (Li et al.,, 2022) and SyneTune (Salinas et al.,, 2022) are internally simulating the runtime, we cannot simulate optimizers of interest if the optimizers are not introduced in the packages. This restricts researchers in simulating new methods, hindering experimentation and fair comparison. Furthermore, their simulation backend assumes that optimizers take the ask-and-tell interface and it requires the reimplementation of optimizers of interest in their codebase. Since reimplementation is time-consuming and does not guarantee its correctness without tests, it is helpful to have an easy-to-use Python wrapper around existing codes. Note that this work extends previous work (Watanabe, 2023a, ), by adding the handling of optimizers with non-negligible overhead and the empirical verification of the simulation algorithm.
@@ 3  Automatic Waiting Time Scheduling Wrapper
As an objective function may take a random seed and fidelity parameters in practice, we denote a set of the arguments for the $n$-th query as $\bm{a}^{(n)}$. In this section, a job means to allocate the $n$-th queried HP configuration $\bm{x}^{(n)}$ to a free worker and obtain its result $r^{(n)}\coloneqq(f^{(n)},\tau^{(n)})=(f(\bm{x}^{(n)}|\bm{a}^{(n)}),\tau(\bm{x}^{(n)}|\bm{a}^{(n)}))$. Besides that, we denote the $n$-th chronologically ordered result as $r_{n}$. Our wrapper outlined in Algorithm 1 is required to satisfy the following conditions:
@@
Note that an example of the restart of evaluation could be when we evaluate DL model instantiated with HP $\bm{x}$ for $20$ epochs and if we want to then evaluate the same HP configuration $\bm{x}$ for $100$ epochs, we start the training of this model from the $21$st epoch instead of from scratch using the intermediate state. Line 4 checks this condition and Line 5 ensures the intermediate state to restart exists before the evaluation. To achieve these features, we chose to share the required information via the file system and create the following JSON files that map:
@@
We additionally provide an approach that also extends to the ask-and-tell interface by providing a Single-Core Simulator (SCS) for single-core scenarios (details omitted for brevity). While the Multi-Core Simulator (MCS) wraps optimizers running with $P$ cores or $P$ workers, SCS runs only on a single core and simulates a $P$-worker run. Unlike previous work (Watanabe, 2023a, ), Algorithm 1 handles expensive optimizers by checking individual workers’ wait times during the latest sampling measured by $t_{\mathrm{now}}$ in Line 12. However, this check complicates race conditions, making it hard to guarantee the correctness of implementation. For this reason, empirical verification through edge cases is provided in the next section.
@@
4:         $\triangleright$ Cond. 1: The new fidelity input in $\bm{a}^{(N+1)}$ must be higher than that in $\bm{a}$ for restart
5:         $\triangleright$ Cond. 2: The registration of $s^{(N+1)}$ to $\mathcal{S}$ must happen before the sample of $\bm{x}^{(N+1)}$
@@
7:     Query the result: $(f^{(N+1)},\tau^{(N+1)})$
8:     Calibrate runtime for restart: $\tau^{(N+1)}\leftarrow\tau^{(N+1)}-\tau$
@@
11:     $\triangleright$ $k$ is the number of results from the other workers that were appended during the wait
@@
16:$\pi$ (an optimizer policy), get_n_results (a function that returns the number of recognized results by our wrapper. The results include the ones that have not been reported to the optimizer yet.).
@@ 4  Empirical Algorithm Verification on Test Cases
In this section, we verify our algorithm using some edge cases. Throughout this section, we use the number of workers $P=4$. We also note that our wrapper behavior depends only on returned runtime at each iteration in a non-continual setup and it is sufficient to consider only runtime $\tau^{(n)}$ and sampling time $t^{(n)}$ at each iteration. Therefore, we use a so-called fixed-configuration sampler, which defines a sequence of HP configurations and their corresponding runtimes at the beginning and samples from the fixed sequence iteratively. More formally, assume we would like to evaluate $N_{\mathrm{all}}$ HP configurations, then the sampler first generates $\{\tau^{(n)}\}_{n=1}^{N_{\mathrm{all}}}$ and one of the free workers receives an HP configuration at the $n$-th sampling that leads to the runtime of $\tau^{(n)}$. Furthermore, we use two different optimizers to simulate the sampling cost:
@@
    Expensive Optimizer: that sleeps for $c(|\mathcal{D}|+1)$ seconds as a sampling overhead before giving $\tau^{(n)}$ to a worker where $|\mathcal{D}|$ is the size of a set of observations and $c\in\mathbb{R}_{+}$ is a proportionality constant, and
@@
    Cheap Optimizer: that gives $\tau^{(n)}$ to a worker immediately without a sampling overhead.
@@ 4.1  Quantitative Verification on Random Test Cases
(c) Cheap optimizer
(d) Expensive optimizer with $c=5\times 10^{-2}$
Figure 3: The return order verification results. When we use our wrapper, the red dots are obtained. If all the dots are aligned on $y=x$, it implies that the return order in a simulation with our wrapper and that in its naïve simulation perfectly match. As expected, the red dots completely overlap with $y=x$. See the text in “Checking Return Orders” for the plot details.
We test our algorithm quantitatively using some test cases. The test cases $\{\tau^{(n)}\}_{n=1}^{N_{\mathrm{all}}}$ where $N_{\mathrm{all}}=100$ for this verification were generated from the following distributions: 1. Uniform $\frac{T}{b}\sim\mathcal{U}(0,2)$, 2. Exponential $\frac{T}{b}\sim\mathrm{Exp}(1)$, 3. Pareto $\frac{T+1}{b}\sim\mathcal{P}(\alpha=1)$, and 4. LogNormal $\ln\frac{\sqrt{e}T}{b}\sim\mathcal{N}(0,1)$, where $T$ is the probability variable of the runtime $\tau$ and we used $b=5$. Each distribution uses the default setups of numpy.random and the constant number $b$ calibrates the expectation of each distribution except for the Pareto distribution to be $5$. Furthermore, we used the cheap optimizer and the expensive optimizer with $c=5\times 10^{-2}$. As $N_{\mathrm{all}}=100$, the worst sampling duration for an expensive optimizer will be $5$ seconds. As we can expect longer waiting times for the expensive optimizer, it is more challenging to yield the precise return order and the precise simulated runtime. Hence, these test cases empirically verify our implementation if our wrapper passes every test case.
@@
We performed the following procedures to check whether the obtained return orders are correct: (1) run optimizations with the naïve simulation (NS), i.e. Figure 1 (Left) and without our wrapper, i.e. Figure 1 (Right), (2) define the trajectories for each optimization $\{\tau^{\mathrm{NS}}_{n}\}_{n=1}^{N_{\mathrm{all}}}$ and $\{\tau_{n}\}_{n=1}^{N_{\mathrm{all}}}$, (3) sort $\{\tau_{n}\}_{n=1}^{N_{\mathrm{all}}}$ so that $\{\tau_{i_{n}}\}_{n=1}^{N_{\mathrm{all}}}=\{\tau^{\mathrm{NS}}_{n}\}_{n=1}^{N_{\mathrm{all}}}$ holds, and (4) plot $\{(n,i_{n})\}_{n=1}^{N_{\mathrm{all}}}$ (see Figure 3). If the simulated return order is correct, the plot $\{(n,i_{n})\}_{n=1}^{N_{\mathrm{all}}}$ will look like $y=x$, i.e. $(n,n)$ for all $n\in[{N_{\mathrm{all}}}]$, and we expect to have such plots for all the experiments. For comparison, we also collect $\{\tau_{n}\}_{n=1}^{N_{\mathrm{all}}}$ without our wrapper, i.e. Figure 1 (Left) without time.sleep in Line 4. As seen in Figure 3, our wrapper successfully replicates the results obtained by the naïve simulation. The test cases by the Pareto distribution are edge cases because it has a heavy tail and it sometimes generates configurations with very long runtime, leading to blue dots located slightly above the red dots. Although this completely confuses the implementation without our wrapper, our wrapper appropriately handles the edge cases.
@@
We check whether the simulated runtimes at each iteration were correctly calculated using the same setups. Figure 4 presents the simulated runtimes for each setup. As can be seen in the figures, our wrapper got a relative error of $1.0\times 10^{-5}\sim 1.0\times 10^{-3}$. Since the expectation of runtime is $5$ seconds except for the Pareto distribution, the error was approximately $0.05\sim 5$ milliseconds and this value comes from the query overhead in our wrapper before each sampling. Although the error is sufficiently small, the relative error becomes much smaller when we use more expensive benchmarks that will give a large runtime $\tau^{(n)}$.
@@ 4.2  Performance Verification on Actual Runtime Reduction
In the previous sections, we verified the correctness of our algorithms and empirically validated our algorithms. In this section, we demonstrate the runtime reduction effect achieved by our wrapper. To test the runtime reduction, we optimized the multi-fidelity 6D Hartmann function 444 We set the runtime function so that the maximum runtime for one evaluation becomes 1 hour. More precisely, we used $10\times r(\bm{z})$ instead of $r(\bm{z})$ in Appendix A.2.  (Kandasamy et al.,, 2017) using random search with $P=4$ workers over 10 different random seeds. In the noisy case, we added a random noise to the objective function. We used both MCS and SCS in this experiment and the naïve simulation. Figure 5 (Left) shows that both MCS and SCS perfectly reproduced the results by the naïve simulation while they finished the experiments $6.7\times 10^{3}$ times and $1.3\times 10^{5}$ times faster, respectively. Note that it is hard to see, but the rightmost curve of Figure 5 (Left) has the three lines: (1) Simulated Runtime (MCS), (2) Simulated Runtime (SCS), and (3) Actual Runtime (Naïve), and they completely overlap with each other. SCS is much quicker than MCS because it does not require communication between each worker via the file system. Although MCS could reproduce the results by the naïve simulation even for the noisy case, SCS failed to reproduce the results because the naïve simulation relies on multi-core optimization, while SCS does not use multi-core optimization. This difference affects the random seed effect on the optimizations. However, since SCS still reproduces the results for the deterministic case, it verifies our implementation of SCS. From the results, we can conclude that while SCS is generally quicker because it does not require communication via the file system, it may fail to reproduce the random seed effect. This is because SCS wraps an optimizer by relying on the ask-and-tell interface instead of using the multi-core implementation provided by the optimizer.